is there Rectangle with 4 square cm and perimeter 10
Answer by AnlytcPhil(1806)

            LW = 4
       2L + 2W = 10

Divide the second equation by 2

         L + W = 5

Solve for L

             L = 5 - W

Substitute in

            LW = 4

      (5 - W)W = 4

       5W - W² = 4

  -W² + 5W - 4 = 0 

Multiply through by -1

   W² - 5W + 4 = 0

(W - 4)(W - 1) = 0

W - 4 = 0;  W - 1 = 0

    W = 4;      W = 1

L = 5 - W       L = 5 - W

L = 5 - 4       L = 5 - 1

L = 1           L = 4

So either a 1cm x 4cm or a 4cm x 1 cm rectangle.

Either of these, which are draw to scale:



Edwin